# Learning under Distributional Drift: Reproducibility as an Intrinsic Statistical Resource

### v1

@@ 1 Introduction
Modern learning systems often operate in self-modifying environments, where learning alters the very distribution it learns from. A recommender policy reshapes user preferences as it updates (Perdomo2020); adaptive experiments change the distribution from which subsequent data are drawn (li2024); and reinforcement-driven agents modify the state transitions that determine future feedback (Sutton1998; Dar2003; AstromMurray2008). In all such cases, the learner is not a passive observer of a stationary process but an active participant in an evolving distributional flow. This feedback coupling breaks not only stationarity but the assumption of statistical reproducibility—that future data could, in principle, be redrawn from the same law (Vapnik1998).
@@ 1.1 Motivation
Given samples $\{(x_{i},y_{i})\}_{i=1}^{T}$ drawn i.i.d. from a fixed distribution $p(x,y)$, population and empirical risks coincide asymptotically, with convergence rate $O(T^{-1/2})$ (Vapnik1998). Once the data-generating process itself changes, these guarantees collapse. When distributional shift depends on the learner’s trajectory, the change is no longer exogenous but endogenous: a function of the learner’s own actions and their cumulative impact on the environment. In such settings, convergence is governed not by sample size, but by how far the underlying distribution moves over the course of learning, a quantity naturally measured in the statistical geometry of the model.
@@ 1.2 Conceptual View
We view the data-generating process as a trajectory $\{\theta_{t}\}$ on a statistical manifold, whose discrete path length represents cumulative information change (AmariNagaoka2000). Any smooth divergence admits a second-order expansion whose Hessian defines a local Riemannian metric (Amari2016). For the broad class of $f$-divergences, including KL, $\chi^{2}$, Hellinger, and $\alpha$-divergences, this induced metric coincides with the Fisher–Rao geometry. Fisher–Rao is therefore the canonical invariant Riemannian metric on a statistical model (cencov1982; AmariNagaoka2000). It is preserved under smooth reparameterizations and captures displacement intrinsic to the model family rather than to any particular coordinate system. Although different divergences quantify global discrepancies between distributions, their local quadratic structure collapses to the same information metric, and only the Fisher metric enjoys this invariance property (cencov1982). At an infinitesimal scale, Fisher information provides the quadratic form that linearizes KL divergence and characterizes the distinguishability of nearby models. This is not a peculiarity of KL, but a reflection of the shared second-order geometry of all $f$-divergences. Consequently, Fisher–Rao arc length offers a natural intrinsic scale for describing how small distributional perturbations alter the statistical environment of a learner.
Accumulating these infinitesimal displacements yields an intrinsic Riemannian arc length (AmariNagaoka2000). This differs fundamentally from summing TV, Wasserstein, or KL increments, which are extrinsic divergences and do not compose additively along a path. What the arc length measures is not merely “how far the distribution moved” in a chosen divergence, but the total statistical distinguishability expended along the trajectory in the unique geometry that governs estimation. Generalization error then depends not only on the number of samples but also on the total path length traversed by the data distribution during learning.
@@ 1.3 Relation to Prior Frameworks
Stationary learning (Vapnik1998; shalev2014) assumes fixed $p(x,y)$ and produces the familiar $O(T^{-1/2})$ convergence guarantees derived from concentration under i.i.d. sampling. Nonstationary stochastic optimization (Besbes2015; Zhao2024) introduces a variation budget that measures how much the target distribution can change exogenously over time. The metric of variation is typically a total-variation or Wasserstein norm, which constrains external motion but does not model coupling to the learner’s policy. Related approaches bound performance under drifting distributions using stability or Rademacher techniques (Blanchard2021; kuznetsov16; Mohri2012) but likewise do not explicitly model feedback between learner and environment. Performative prediction (Perdomo2020) does explicitly introduce feedback between model and environment, focusing on equilibrium conditions where that feedback vanishes ($p_{\theta_{t+1}}\!\approx\!p_{\theta_{t}}$). Follow-up work on performative prediction has refined aspects of this coupling (izzo21; jagadeesan22), but continues to analyze convergence to fixed points rather than the explicit accumulation of temporal drift. Adaptive data analysis (Dwork2015; rogers20; bassily2016) addresses feedback in the statistical query setting and bounds overfitting through mutual-information or stability arguments, yet does so within a static population that does not itself evolve (see also russo2016; Seldin2012; AlquierGuedj2018).
@@
Stationary i.i.d. $\sqrt{T}$ term only Vapnik1998 No drift or feedback; fixed $p(x,y)$.
Exogenous drift $d_{t}$ term Besbes2015 Distribution changes externally over time.
Performative equilibrium $d_{t}\!\to\!0,\ \kappa_{t}^{(\mathcal{M})}\!\to\!0$ Perdomo2020 Coupled learner–environment system converging to a performative equilibrium.
Adaptive data analysis $\kappa_{t}^{(\mathcal{M})}$ term Dwork2015 Purely feedback-driven nonreproducibility.
A broader literature also bounds performance in nonstationary and online optimization through dynamic regret or excess-risk budgets (Zhao2024), while information-theoretic analyses of adaptive learning express related limits in terms of mutual-information accumulation between data and model parameters (russo2016; xu2017). In statistical physics, learning has been interpreted as a thermodynamic process in which information gain is constrained by entropy production (Still2012; Goldt2017; ortega2013thermodynamics). All of these perspectives describe some notion of limited information flow, but they do so with external metrics or equilibrium arguments rather than an intrinsic geometric invariant.
@@ 2.2 Regularity and Curvature Bounds
    and the $O(\|x\|^{2})$ terms are controlled uniformly over $\theta\in K$. These bounds are used only to control the quadratic remainder in Lemma 4.
@@
Let $(\Theta,g)$ be a statistical manifold equipped with the Fisher metric. Fix $\theta\in\Theta$ and suppose the sectional curvature is bounded in a neighborhood of $\theta$. Then there exists $r>0$ and a constant $C>0$ such that for all $\Delta\theta$ with $\|\Delta\theta\|_{g_{\theta}}\leq r$,
$d_{F}(\theta,\theta+\Delta\theta)=\|\Delta\theta\|_{g_{\theta}}\,\bigl(1+E(\Delta\theta)\bigr),\qquad|E(\Delta\theta)|\leq C\,\|\Delta\theta\|_{g_{\theta}}.$ (5)
@@
Proof [Proof sketch] In normal coordinates at $\theta$, the Fisher metric satisfies $g_{ij}(x)=\delta_{ij}+O(\|x-\theta\|^{2})$, and the exponential map obeys $\operatorname{Exp}_{\theta}(v)=\theta+v+O(\|v\|^{2})$ (AmariNagaoka2000; Lee97). Thus if $\theta^{\prime}=\theta+\Delta\theta$ is sufficiently close to $\theta$, inverting the latter relation yields $v=\Delta\theta+O(\|\Delta\theta\|^{2})$. Gauss’s lemma then gives $d_{F}(\theta,\theta^{\prime})=\|v\|_{g_{\theta}}=\|\Delta\theta\|_{g_{\theta}}+O(\|\Delta\theta\|_{g_{\theta}}^{2})$, and shrinking the neighborhood of $\theta$ yields the stated uniform two–sided inequality.
@@ 2.3 Concentration Tools
This is a standard sub-Gaussian martingale concentration argument (Boucheron2013; bercu2015).
@@ 2.4 Length Convergence and Path Integrals
2.5 Summary of Constants
For reference, the principal constants are:
$\sigma$    sub-Gaussian loss parameter
$L_{p}$ Lipschitz constant linking risk to $W_{1}$
$G_{\max}$ bound on Fisher curvature (Christoffel symbols)
$c_{d}$ local geodesic equivalence constant
@@ 3 Problem Setup
where the matrices $A$ and $B$ determine sensitivity to the learner’s action $u_{t}$ and to exogenous influences $\eta_{t}$ (AstromMurray2008). The term $Au_{t}$ captures policy-induced motion and vanishes in the purely exogenous case $A=0$, while the term $B\eta_{t}$ captures background evolution and vanishes when $B=0$, yielding drift driven entirely by the learner.
@@
A natural question is whether such coupling might guide the process toward a performative equilibrium, so that policy-induced drift $\kappa_{t}^{(\mathcal{M})}$ eventually vanishes. In general, this need not occur. The joint evolution of $(f_{t},\theta_{t})$ is not assumed to be a gradient flow on a shared potential, and the environment map $F$ need not be contractive. Even when fixed points exist, stochasticity, misspecification, or misaligned objectives can yield persistent cycles or wandering trajectories. Prior work on performative prediction typically imposes strong stability conditions to study equilibria (Perdomo2020). Our analysis instead addresses the generic case in which the learner–environment system continues to move along the statistical manifold and the cumulative motion grows with $T$.
@@
a tractable surrogate for the Fisher path length of the trajectory. The scalar $\alpha>0$ is not a model parameter. It plays no role in the dynamics of $\theta_{t}$ and serves only to place $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ on a common scale so they can be combined into the single quantity $C_{T}$.
@@ 4.1 Drift Decomposition
where $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ are the drift magnitudes from Section 3, $\alpha$ balances their relative scale, and $\epsilon_{t}$ consists of second–order residuals with $\mathbb{E}[\epsilon_{t}]\penalty 10000\ =\penalty 10000\ O\!\left(\mathbb{E}[\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2}]\right)$, under the smoothness conditions on $F$ (AmariNagaoka2000; Absil2008). Equation (10) gives the basic link between first–order motion and its geometric decomposition. While the decomposition is algebraically simple, its validity in the intrinsic Fisher–Rao geometry is not automatic. It relies on the smoothness and curvature controls established in Section 2.
@@ 5.3 Drift–Feedback Generalization Bound
This follows by combining the drift–feedback bound (Theorem 13) with standard PAC–Bayes inequalities. See, for example, Seldin2012; AlquierGuedj2018.
@@ 5.4 Lower Bound and Speed Limit
Choosing $\delta$ small enough ensures $m\delta^{2}\leq\alpha\log|V|$ for some $\alpha<1$, which is the Fano separation condition. Because all trajectories remain inside an $O(\delta)$ neighborhood of $\theta_{0}$, Assumption 1 ensures uniform validity of the quadratic KL expansion and bounded curvature. The full technical justification is given in Appendix A.4.
The multi–hypothesis Fano bound then yields a minimax error of order $C/T$ over all processes with $C_{T}\leq C$. When $C_{T}=0$, the same argument recovers the classical stationary $\Omega(T^{-1/2})$ lower bound, giving the combined $\max(T^{-1/2},\,C/T)$ rate.
@@ 5.5 Stationary and Classical Limits
  • i.i.d. regime. If $d_{t}\equiv 0$ and $J_{u}F\equiv 0$, then $\mathbb{E}|\hat{R}_{T}-R_{T}|=O(T^{-1/2})$, matching classical generalization results (Vapnik1998).
  • Exogenous drift. If $J_{u}F\equiv 0$ but $d_{t}>0$, the term $(C_{1}/T)\sum_{t}d_{t}$ reduces to the variation–budget penalty of Besbes2015.
  • Performative equilibrium. If both drift components vanish asymptotically, the bound returns to $O(T^{-1/2})$ (Perdomo2020).
  • Adaptive data analysis. If $d_{t}\equiv 0$ but $\kappa_{t}^{(\mathcal{M})}>0$, the $(C_{2}/T)\sum_{t}\kappa_{t}^{(\mathcal{M})}$ term parallels stability penalties in adaptive data analysis (Dwork2015).
@@ 6 Geometric and Information–Theoretic Interpretation
A brief remark clarifies why Fisher–Rao geometry is a natural scale on which to measure statistical motion. Many common distances between distributions—such as total variation, Wasserstein distance, or mutual-information measures—are defined directly on probability laws and therefore do not reflect the geometric structure of a parametric model family. Fisher–Rao, in contrast, is the unique Riemannian metric on a statistical manifold that is invariant under smooth reparameterizations of the model (cencov1982), and so captures displacement intrinsic to the family rather than to any particular coordinate system. Moreover, any smooth divergence admits a second-order expansion that induces a local Riemannian metric (Amari2016). For the broad class of $f$-divergences (including KL, reverse-KL, $\chi^{2}$, Hellinger, and $\alpha$-divergences) this metric coincides with Fisher–Rao. This distinction matters in the endogenous setting, where the learner’s action enters through $J_{u}F$ and the statistical effect of a control depends on the local curvature of the model class. Fisher–Rao arc length therefore provides a curvature-aware, coordinate-invariant notion of statistical motion and supplies an appropriate local scale for describing drift on the model manifold.
@@
then reflects the fact that no learner can achieve smaller worst–case generalization error than that imposed by the cumulative Fisher–Rao motion of the data–generating distribution. Any process that accumulates substantial distributional motion necessarily incurs a proportionally large generalization gap. Classical quantities used to measure nonstationarity, such as Wasserstein variation budgets and mutual-information constraints, correspond to coordinate projections of the same second–order structure. The Fisher–Rao arc length $C_{T}$ records its intrinsic form and is therefore the appropriate quantity controlling reproducibility under self-altering dynamics. In this sense, $C_{T}$ is mathematically analogous to an accumulated information–production or entropy–production functional (Still2012; Goldt2017; ortega2013thermodynamics), aggregating the irreversibility induced by both exogenous changes and policy sensitivity.

### v2

@@ 1 Introduction
Modern learning systems often operate in self-modifying environments, where learning alters the very distribution it learns from. A recommender policy reshapes user preferences as it updates (Perdomo et al., 2020); adaptive experiments change the distribution from which subsequent data are drawn (Li et al., 2024); and reinforcement-driven agents modify the state transitions that determine future feedback (Sutton and Barto, 2018; Even‐Dar and Mansour, 2003). In all such cases, the learner is not a passive observer of a stationary process but an active participant in an evolving distributional flow. This feedback coupling breaks not only stationarity but the assumption of statistical reproducibility—that future data could, in principle, be redrawn from the same law (Vapnik, 1998).
@@ 1.1 Motivation
Given samples $\{(x_{i},y_{i})\}_{i=1}^{T}$ drawn i.i.d. from a fixed distribution $p(x,y)$, population and empirical risks coincide asymptotically, with convergence rate $O(T^{-1/2})$ (Vapnik, 1998). Once the data-generating process itself changes, these guarantees collapse. When distributional shift depends on the learner’s trajectory, the change is no longer exogenous but endogenous: a function of the learner’s own actions and their cumulative impact on the environment. In such settings, convergence is governed not by sample size, but by how far the underlying distribution moves over the course of learning, a quantity naturally measured in the statistical geometry of the model.
@@ 1.2 Conceptual View
We view the data-generating process as a trajectory $\{\theta_{t}\}$ on a statistical manifold, whose discrete path length represents cumulative information change (Amari and Nagaoka, 2000). Any smooth divergence admits a second-order expansion whose Hessian defines a local Riemannian metric (Amari, 2016). For the broad class of $f$-divergences, including KL, $\chi^{2}$, Hellinger, and $\alpha$-divergences, this induced metric coincides with the Fisher–Rao geometry. Fisher–Rao is therefore the canonical invariant Riemannian metric on a statistical model (Čencov, 1982; Amari and Nagaoka, 2000). It is preserved under smooth reparameterizations and captures displacement intrinsic to the model family rather than to any particular coordinate system. Although different divergences quantify global discrepancies between distributions, their local quadratic structure collapses to the same information metric, and only the Fisher metric enjoys this invariance property (Čencov, 1982). At an infinitesimal scale, Fisher information provides the quadratic form that linearizes KL divergence and characterizes the distinguishability of nearby models. This is not a peculiarity of KL, but a reflection of the shared second-order geometry of all $f$-divergences. Consequently, Fisher–Rao arc length offers a natural intrinsic scale for describing how small distributional perturbations alter the statistical environment of a learner.
Accumulating these infinitesimal displacements yields an intrinsic Riemannian arc length (Amari and Nagaoka, 2000). This differs fundamentally from summing TV, Wasserstein, or KL increments, which are extrinsic divergences and do not compose additively along a path. What the arc length measures is not merely “how far the distribution moved” in a chosen divergence, but the total statistical distinguishability expended along the trajectory in the unique geometry that governs estimation. Generalization error then depends not only on the number of samples but also on the total path length traversed by the data distribution during learning.
@@ 1.3 Relation to Prior Frameworks
Stationary learning (Vapnik, 1998; Shalev-Shwartz and Ben-David, 2014) assumes fixed $p(x,y)$ and produces the familiar $O(T^{-1/2})$ convergence guarantees derived from concentration under i.i.d. sampling. Nonstationary stochastic optimization (Besbes et al., 2015; Zhao et al., 2024) introduces a variation budget that measures how much the target distribution can change exogenously over time. The metric of variation is typically a total-variation or Wasserstein norm, which constrains external motion but does not model coupling to the learner’s policy. Related approaches bound performance under drifting distributions using stability or Rademacher techniques (Blanchard et al., 2021; Kuznetsov and Mohri, 2016; Mohri and Muñoz Medina, 2012) but likewise do not explicitly model feedback between learner and environment. PAC–Bayes analyses have also been developed for distributional change, notably in the domain adaptation setting of Germain et al. (2016). However, those results bound a single source–target shift, whereas the present work studies path-dependent, cumulative drift induced by sequential learner–environment coupling. Performative prediction (Perdomo et al., 2020) does explicitly introduce feedback between model and environment, focusing on equilibrium conditions where that feedback vanishes ($p_{\theta_{t+1}}\!\approx\!p_{\theta_{t}}$). Follow-up work on performative prediction has refined aspects of this coupling (Izzo et al., 2021; Jagadeesan et al., 2022), but continues to analyze convergence to fixed points rather than the explicit accumulation of temporal drift. Adaptive data analysis (Dwork et al., 2015; Rogers et al., 2020; Russo and Zou, 2016; Bassily et al., 2016) addresses feedback in the statistical query setting and bounds overfitting through mutual-information or stability arguments, yet does so within a static population that does not itself evolve.
@@
Stationary i.i.d. $\sqrt{T}$ term only Vapnik (1998) No drift or feedback; fixed $p(x,y)$.
Exogenous drift $d_{t}$ term Besbes et al. (2015) Distribution changes externally over time.
Performative equilibrium $d_{t}\!\to\!0,\ \kappa_{t}^{(\mathcal{M})}\!\to\!0$ Perdomo et al. (2020) Coupled learner–environment system converging to a performative equilibrium.
Adaptive data analysis $\kappa_{t}^{(\mathcal{M})}$ term Dwork et al. (2015) Purely feedback-driven nonreproducibility.
A broader literature also bounds performance in nonstationary and online optimization through dynamic regret or excess-risk budgets (Zhao et al., 2024), while information-theoretic analyses of adaptive learning express related limits in terms of mutual-information accumulation between data and model parameters (Russo and Zou, 2016; Xu and Raginsky, 2017). In statistical physics, learning has been interpreted as a thermodynamic process in which information gain is constrained by entropy production (Still et al., 2012; Goldt and Seifert, 2017; Ortega and Braun, 2013). All of these perspectives describe some notion of limited information flow, but they do so with external metrics or equilibrium arguments rather than an intrinsic geometric invariant.
@@ 2.2 Regularity and Curvature Bounds
    and the $O(\|x\|^{2})$ terms are controlled uniformly over $\theta\in K$. These bounds ensure the existence of a uniform normal neighborhood around each $\theta\in K$, within which Lemma 4 holds.
@@
Let $(\Theta,g)$ be a statistical manifold equipped with the Fisher metric. Fix $\theta\in\Theta$ and suppose the sectional curvature is bounded in a neighborhood of $\theta$. Throughout, expressions of the form $\theta+\Delta\theta$ are understood in normal coordinates centered at $\theta$. Then there exists $r>0$ and a constant $C>0$ such that for all $\Delta\theta$ with $\|\Delta\theta\|_{g_{\theta}}\leq r$,
$d_{F}(\theta,\theta+\Delta\theta)=\|\Delta\theta\|_{g_{\theta}}.$ (5)
@@
Proof [Proof sketch] Work in normal coordinates centered at $\theta$. By definition of these coordinates, for $\|\Delta\theta\|_{g_{\theta}}$ sufficiently small the point $\theta+\Delta\theta$ lies in a normal neighborhood and satisfies $\theta+\Delta\theta=\operatorname{Exp}_{\theta}(\Delta\theta)$; see Lee (1997, Prop. 5.11). Gauss’s lemma then gives
$d_{F}(\theta,\theta+\Delta\theta)=\|\Delta\theta\|_{g_{\theta}}.$
The stated two–sided inequality follows immediately from this identity for any $C>0$ after restricting to $\|\Delta\theta\|_{g_{\theta}}\leq r$.
@@ 2.3 Concentration Tools
A proof is provided in A.
@@ 3 Problem Setup
where the matrices $A$ and $B$ determine sensitivity to the learner’s action $u_{t}$ and to exogenous influences $\eta_{t}$ (Åström and Murray, 2008). The term $Au_{t}$ captures policy-induced motion and vanishes in the purely exogenous case $A=0$, while the term $B\eta_{t}$ captures background evolution and vanishes when $B=0$, yielding drift driven entirely by the learner.
@@
A natural question is whether such coupling might guide the process toward a performative equilibrium, so that policy-induced drift $\kappa_{t}^{(\mathcal{M})}$ eventually vanishes. In general, this need not occur. The joint evolution of $(f_{t},\theta_{t})$ is not assumed to be a gradient flow on a shared potential, and the environment map $F$ need not be contractive. Even when fixed points exist, stochasticity, misspecification, or misaligned objectives can yield persistent cycles or wandering trajectories. Prior work on performative prediction typically imposes strong stability conditions to study equilibria (Perdomo et al., 2020). Our analysis instead addresses the generic case in which the learner–environment system continues to move along the statistical manifold and the cumulative motion grows with $T$.
@@
a tractable surrogate for the Fisher path length of the trajectory. The scalar $\alpha>0$ is not a model parameter. It is a fixed norm-equivalence factor that absorbs local scaling differences between exogenous drift and control-induced tangent components under the Fisher metric, so they can be combined into the single quantity $C_{T}$.
@@ 4.1 Drift Decomposition
where $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$ are the drift magnitudes from Section 3, $\alpha$ balances their relative scale, and $\epsilon_{t}$ consists of second–order residuals with $\mathbb{E}[\epsilon_{t}]\penalty 10000\ =\penalty 10000\ O\!\left(\mathbb{E}[\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2}]\right)$, under the smoothness conditions on $F$ (Amari and Nagaoka, 2000; Lee, 1997). Equation (10) gives the basic link between first–order motion and its geometric decomposition. While the decomposition is algebraically simple, its validity in the intrinsic Fisher–Rao geometry is not automatic. It relies on the smoothness and curvature controls established in Section 2.
@@ 5.3 Drift–Feedback Generalization Bound
The bound is obtained by applying the PAC–Bayes Hoeffding–Azuma inequality for martingales (Seldin et al., 2012, Theorem 6), which extends the classical PAC–Bayes framework of McAllester (1998), to the stochastic term remaining after the drift–feedback decomposition of Theorem 13.
@@ 5.4 Lower Bound and Speed Limit
Choosing $\delta$ small enough ensures $m\delta^{2}\leq\zeta\log|V|$ for some $\zeta<1$, which is the Fano separation condition. Because all trajectories remain inside an $O(\delta)$ neighborhood of $\theta_{0}$, Assumption 1 ensures uniform validity of the quadratic KL expansion and bounded curvature. The full technical justification is given in Appendix A.4.
By Fano’s inequality (Yu, 1997), this yields a minimax error of order $C/T$ over all processes with $C_{T}\leq C$. When $C_{T}=0$, the same argument recovers the classical stationary $\Omega(T^{-1/2})$ lower bound, giving the combined $\max(T^{-1/2},\,C/T)$ rate.
@@ 5.5 Stationary and Classical Limits
  • i.i.d. regime. If $d_{t}\equiv 0$ and $J_{u}F\equiv 0$, then $\mathbb{E}|\hat{R}_{T}-R_{T}|=O(T^{-1/2})$, matching classical generalization results (Vapnik, 1998).
  • Exogenous drift. If $J_{u}F\equiv 0$ but $d_{t}>0$, the term $(C_{1}/T)\sum_{t}d_{t}$ is analogous to the variation–budget penalty of Besbes et al. (2015).
  • Performative equilibrium. If both drift components vanish asymptotically, the bound returns to $O(T^{-1/2})$ (Perdomo et al., 2020).
  • Adaptive data analysis. If $d_{t}\equiv 0$ but $\kappa_{t}^{(\mathcal{M})}>0$, the $(C_{2}/T)\sum_{t}\kappa_{t}^{(\mathcal{M})}$ term parallels stability penalties in adaptive data analysis (Dwork et al., 2015; Russo and Zou, 2016; Bassily et al., 2016).
@@ 6 Geometric and Information–Theoretic Interpretation
A brief remark clarifies why Fisher–Rao geometry is a natural scale on which to measure statistical motion. Many common distances between distributions—such as total variation, Wasserstein distance, or mutual-information measures—are defined directly on probability laws and therefore do not reflect the geometric structure of a parametric model family. Fisher–Rao, in contrast, is the unique Riemannian metric on a statistical manifold that is invariant under smooth reparameterizations of the model (Čencov, 1982), and so captures displacement intrinsic to the family rather than to any particular coordinate system. Moreover, any smooth divergence admits a second-order expansion that induces a local Riemannian metric (Amari, 2016). For the broad class of $f$-divergences (including KL, reverse-KL, $\chi^{2}$, Hellinger, and $\alpha$-divergences) this metric coincides with Fisher–Rao. This distinction matters in the endogenous setting, where the learner’s action enters through $J_{u}F$ and the statistical effect of a control depends on the local curvature of the model class. Fisher–Rao arc length therefore provides a curvature-aware, coordinate-invariant notion of statistical motion and supplies an appropriate local scale for describing drift on the model manifold.
@@
then reflects the fact that no learner can achieve smaller worst–case generalization error than that imposed by the cumulative Fisher–Rao motion of the data–generating distribution. Any process that accumulates substantial distributional motion necessarily incurs a proportionally large generalization gap. Classical quantities used to measure nonstationarity, such as Wasserstein variation budgets and mutual-information constraints, correspond to coordinate projections of the same second–order structure. The Fisher–Rao arc length $C_{T}$ records its intrinsic form and is therefore the appropriate quantity controlling reproducibility under self-altering dynamics. In this sense, $C_{T}$ is mathematically analogous to an accumulated information–production or entropy–production functional (Still et al., 2012; Goldt and Seifert, 2017; Ortega and Braun, 2013), aggregating the irreversibility induced by both exogenous changes and policy sensitivity.